Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Q.

Expert replies
by nikvana » Wed Dec 03, 2008 1:48 am
Kurt, a painter, has 9 jars of paint:
4 are yellow
2 are red
rest are brown
Kurt will combine 3 jars of paint into a new container to make a new color, which he will name accordingly to the following conditions:

Brun Y if the paint contains 2 jars of brown paint and no yellow
Brun X if the paint contains 3 jars of brown paint
Jaune X if the paint contains at least 2 jars of yellow
Jaune Y if the paint contains exactly 1 jar of yellow

What is the probability that the new color will be Jaune

a) 5/42
b) 37/42
c) 1/21
d) 4/9
e) 5/9

Sol: 1. This has at least 2 yellow meaning..

a> there can be all three Y => 4c3
OR
b> 2 Y and 1 out of 2 R and 3 B => 4c2 x 5c1

Total 34

2.This has exactly 1 Y and remaining 2 out of 5 = > 4c1 x 5c2

Total 40
Total possibilities = (9!/3!6!) = 84
Adding the two probabilities: probability = 74/84 = 37/42


I solved it like
prob of it being Jaune X = 4/9 x 3/8 x 2/7 + 4/9 x 3/8 x 5/7 = 1/3
prob of it being Jaune Y = 4/9 x 5/8 x 4/7 = 10/63

Total prob of being jaune = 1/3 + 10/63 = 31/63.
Whats wrong with this method? :?:
Join the discussion
Source: — Problem Solving |

by tanviet » Wed Dec 03, 2008 5:39 am
I do not understand the question. there is only 1 brown jar. How can we have 2 brown jar in the conditions.
Join the discussion

by xyzabc123 » Wed Dec 03, 2008 6:04 am
I did it differently.
if P (Jaune X/Y) needs to be calculated, all combinations containing at least one yellow jar need to be accounted for.
So the opposite will be calculating P(not Jaune X/Y). This corresponds to answering the question: what is the probability that a selection of three jars will have no yellow jar?

5/9 * 4/8 * 3/7 = 5/42

So P(Jaune X/Y) = 1 - 5/42 = 37/41

choose B
Join the discussion

by mental » Wed Dec 03, 2008 6:07 am
I solved it like
prob of it being Jaune X = 4/9 x 3/8 x 2/7 + 4/9 x 3/8 x 5/7 = 1/3
prob of it being Jaune Y = 4/9 x 5/8 x 4/7 = 10/63

Total prob of being jaune = 1/3 + 10/63 = 31/63.
Sol II:
prob of it being Jaune X = (4/9 x 3/8 x 2/7) + (4/9 x 3/8 x 5/7 x 3C1) =
17/42
(3C1 = ways in which non yellow paint can be chosen: YYX, YXY, XYY)

prob of it being Jaune Y = (4/9 x 5/8 x 4/7 x 3C1) = 20/42
(3C1 = ways in which yellow paint can be chosen: xxY, xYx, Yxx)

17/42 + 20/42 = 37/42

..................................................................[/quote]
Join the discussion

by mental » Wed Dec 03, 2008 6:09 am
I solved it like
prob of it being Jaune X = 4/9 x 3/8 x 2/7 + 4/9 x 3/8 x 5/7 = 1/3
prob of it being Jaune Y = 4/9 x 5/8 x 4/7 = 10/63

Total prob of being jaune = 1/3 + 10/63 = 31/63.
Sol II:
prob of it being Jaune X = (4/9 x 3/8 x 2/7) + (4/9 x 3/8 x 5/7 x 3C1) =
17/42
(3C1 = ways in which non yellow paint can be chosen: YYX, YXY, XYY)

prob of it being Jaune Y = (4/9 x 5/8 x 4/7 x 3C1) = 20/42
(3C1 = ways in which yellow paint can be chosen: xxY, xYx, Yxx)

17/42 + 20/42 = 37/42

..................................................................[/quote]
Join the discussion